OpenStudy (anonymous):

N_n is the number of ordered n-tuples (a_1,a_2, ..., a_n) that satisfy 1/a_1 + 1/a_2 + ... + 1/a_n = 1. Is N_10 even or odd?

OpenStudy (kinggeorge):

I'm assuming each \(a_i\) must be an integer correct?

OpenStudy (anonymous):

yes, positive integer

OpenStudy (kinggeorge):

Let's say that \(a_1=2\). Then we have that \[{1 \over a_2}+{1 \over a_3}+...+{1 \over a_{10}}={1 \over 2}\]The simplest way to do this, is each \(a_j=18\) so that we have \[{1 \over 18}+{1 \over 18}+...+{1 \over 18}=9 \cdot {1 \over 18}=3\cdot {1 \over 6}\] So for three \(a_1, a_2, a_3\), we need to find how many ways \[{1 \over a_1}+{1 \over a_2}+{1 \over a_3}={1 \over 6}\]

OpenStudy (anonymous):

I don't see how this will work. Not every combination will include an a_i = 2, nor 3 a_j's whose inverses add to 1/6. You can't find every combination using this method. Anyways, I think the trick is to look at what the answer is asking for: even or odd, so it probably has to do with pairs.

OpenStudy (kinggeorge):

You're probably right. This is too much of a brute force method.

OpenStudy (kinggeorge):

Let's start with small cases then. If \(n=1\), we obviously only have 1 way to do it.

OpenStudy (kinggeorge):

If \(n=2\), we only have the possibility that \(a_1=a_2=2\) So again, one possibility. The next step is where it starts getting a little trickier.

OpenStudy (kinggeorge):

If \(n=3\), we have the cases that \[a_1=a_2=a_3=3\]and that two of those \(a_j\) are 4, and the other is 4. Thus, we have 1+3=4 ways to do this if \(n=3\).

OpenStudy (kinggeorge):

I should have said "and the other is 2."

OpenStudy (kinggeorge):

After that, I'm lost for now.

OpenStudy (anonymous):

For n=3, if we have a solution (a1,a2,a3,...,an), then we can generate another one by switching a1 and a2, so since we are looking for even or odd, any solution with a1=/=a2 can be disregarded, so we can set a1=a2, get 1/a1+1/a2+1/a3=2/a1+1/a3=1 so, a1-a1a3+2a3 = 0 => (a1-2)(a3-1)=-2, so the even or odd-ness of the solutions will depend on that of the number of 2 factor combinations of factors of -2, or -1,2; 1,-2, so N_3 is even. I think that this is extendable to the case with n=10

OpenStudy (anonymous):

Anyways, thanks for the help. Just so you know, this is an old Putnam problem... I think I have it figured out now.

OpenStudy (kinggeorge):

Putnam... That would explain it. What year?

OpenStudy (anonymous):

I don't remember... 1997 maybe? or '79? I think it was problem A3 that year.

OpenStudy (anonymous):

1997 A5

OpenStudy (kinggeorge):

That's an interesting solution I'm getting. I'm not quite following it.

OpenStudy (kinggeorge):

From what I'm understanding, the way I first approached the problem is closest to the best way to do it, but I did it backwards.